OpenStudy (anonymous):

Find the equation of the line that passes through (-2,5) and is perpendicular to 2x+2y=18

OpenStudy (danjs):

recall , perpendicular slopes are a negative reciprocals of each other, m -- -1/m

OpenStudy (anonymous):

so the answer would be 2, -5 then?

OpenStudy (vickyisthesmartone123):

Yeah they are basically already giving you it...

OpenStudy (danjs):

find what the slope of the given line is, then yhou know the perpendicular slope, given a point , now have a slope, you can make the equation for the line

OpenStudy (vickyisthesmartone123):

So do we fill in the x's with the point's and we get our answer?

OpenStudy (anonymous):

I don't know how to do any of that lol

OpenStudy (danjs):

given line 2x + 2y = 18 slope of that , if you remember the m value in the form y = mx+b is the slope, will be y = -x + 9, the slope for that one is -1

OpenStudy (anonymous):

so it would be 8?

OpenStudy (vickyisthesmartone123):

@DanJS OHHH..ok I get it now

OpenStudy (danjs):

so the slope of a line perpendicular to that given line will be -1/-1 or +1

OpenStudy (anonymous):

so is that the answer?

OpenStudy (anonymous):

-1/-1 or +1?

OpenStudy (danjs):

you have the slope, +1 and a point (-2 , 5) given on the line, you can make an equation with those

OpenStudy (danjs):

yeah anything divided by that same thing, is 1

OpenStudy (anonymous):

how? by inputing them into the equation already given?

OpenStudy (danjs):

no you recall this y - y1 = m*(x - x1) the point slope line form, you know the slope m, and the point (x1 , y1) just put those in

OpenStudy (danjs):

y - 5 = 1 * (x - (-2)) simplify to some other from if you want, probably same standard as the other line

OpenStudy (anonymous):

ok so what is the answer? lol

OpenStudy (anonymous):

that's the part I always get stuck on

OpenStudy (anonymous):

I know how to do the work, but I always get it wrong

OpenStudy (vickyisthesmartone123):

-1 is the answer @AlixxKK

OpenStudy (anonymous):

oh lol thanks

OpenStudy (danjs):

they want the equation of a line , perpendicular to that other line, and through that point

OpenStudy (anonymous):

ok